1. Using the Frobenius method solve the following ODE:

x2y" + 4xy' + (x2 + 2)y = 0


2. Determine the first and second order partial derivatives for the function:

u(x,t) = Ce(1-n²π²)tsin(nπx)


3. Obtain the Fourier series for the following function:

0 -π < x < 0

f(x) =

sin x 0 < x < π


1
Expert's answer
2021-11-24T15:46:27-0500

Let the solution of above differential equation be

"y=a_ox^m+a_1x^{m+1}+a_2x^{m+2}+......................+a_nx^{n+m}"


"\dfrac{dy}{dx}=ma_ox^{m-1}+(m+1)a_1x^m+(m+2)a_2x^{m+1}+........................."


"\dfrac{d^2y}{dx^2}=m(m-1)a_ox^{m-2}+(m+1)ma_1x^{m-1}+(m+2)(m+1)a_2x^{m}+.................."



"x^2\dfrac{d^2y}{dx^2}+4x\dfrac{dy}{dx}+(x^2+2)y=0"


"x^2[m(m-1)a_ox^{m-2}+(m+1)ma_1x^{m-1}+(m+2)(m+1)a_2x^{m}+..................]+4x[ma_ox^{m-1}+(m+1)a_1x^m+(m+2)a_2x^{m+1}+.........................]+(x^2+2)[a_ox^m+a_1x^{m+1}+a_2x^{m+2}+......................+a_nx^{n+m}]=0"


Equating the coefficient of lowest degree in x (which is xm) equal to zero

"m(m-1)a_o+4ma_o+2a_o=0"

"m^2+3m+2=0"

"m=-2,-1"


For further powers of x,

"x^{m+1}":

"m(m+1)a_1+4(m+1)a_1+2a_1=0"

"m^2+m+4m+4+2=0"

"m^2+5m+6=0"

"m=-2,-3"


Since the equation has three roots it cannot be solved by frobenius method

The given differential equation refers to Sturm-Liouville equation.


2.

"u(x,t) = Ce^{(1-n\u00b2\u03c0\u00b2)t}sin(n\u03c0x)"


"u_x=n\pi Ce^{(1-n²π²)t}cos(nπx)"


"u_t= C(1-n\u00b2\u03c0\u00b2) e^{(1-n\u00b2\u03c0\u00b2)t}sin(n\u03c0x)"


"u_{xx}=-n^2\pi^2 Ce^{(1-n²π²)t}sin(nπx)"


"u_{tt}= C(1-n\u00b2\u03c0\u00b2)^2 e^{(1-n\u00b2\u03c0\u00b2)t}sin(n\u03c0x)"


"u_{xt}= n\pi C(1-n²π²) e^{(1-n²π²)t}cos(nπx)"


3.

"f(x)=a_0+\displaystyle{\sum_{n=1}^{\infin}}a_ncos(n\pi x/L)+\displaystyle{\sum_{n=1}^{\infin}}b_nsin(n\pi x/L)"

we have: "L=\pi"


"a_0=\frac{1}{2L}\int^L_{-L}f(x)dx=\frac{1}{2\pi}\int^{\pi}_{0}sinxdx=\frac{1}{\pi}"


"a_n=\frac{1}{L}\int^L_{-L}f(x)cos(n\pi x/L)dx=\frac{1}{\pi}\int^{\pi}_{0}sinxcos(nx)dx="


"=\frac{nsinxsin(nx)+cosxcos(nx)}{\pi (n^2-1)}|^{\pi}_0=-\frac{cos(n\pi)+1}{\pi (n^2-1)}"


"b_n=\frac{1}{L}\int^L_{-L}f(x)sin(n\pi x/L)dx=\frac{1}{\pi}\int^{\pi}_{0}sinxsin(nx)dx="


"=\frac{cosxsin(nx)-nsinxcos(nx)}{\pi (n^2-1)}|^{\pi}_0=0"


"f(x)=\frac{1}{\pi}-\displaystyle{\sum_{n=1}^{\infin}}\frac{cos(n\pi)+1}{\pi (n^2-1)}cos(n x)"